40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 21 15:21:07 2013, yuta, Update, VAC, all heavy doors on, started pumping down annulus 
    Reply  Thu Feb 21 16:53:37 2013, Manasa, Update, PSL, PSL shutter 
       Reply  Fri Feb 22 13:03:49 2013, Manasa, Update, , MC REFL PD murder dead.jpg
          Reply  Fri Feb 22 17:33:25 2013, Manasa, Update, Electronics, MC REFL PD back from the dead 
             Reply  Wed Apr 22 17:04:11 2015, rana, Update, Electronics, MC REFL PD back from the dead 
                Reply  Thu Apr 23 08:43:40 2015, Steve, Update, Electronics, EG&G InGaAs diodes in stock 
    Reply  Thu Feb 21 18:01:27 2013, Steve, Update, VAC, pumpdown at 230 Torr 
       Reply  Thu Feb 21 21:12:42 2013, Koji, Update, VAC, pumpdown at 230 Torr 
          Reply  Fri Feb 22 08:01:25 2013, Steve, Update, VAC, pumpdown continuous 16hrs_pd75.png
             Reply  Fri Feb 22 11:12:16 2013, Steve, Update, VAC, pumpdown completed mag60min.pngpd75at20h.png
                Reply  Mon Feb 25 08:22:03 2013, Steve, Update, VAC, pumpdown at day 4 4d.png66d@atm.png
                   Reply  Fri Mar 1 14:50:36 2013, Steve, Update, VAC, pumpdown day 8 bgTP3d7.pngRGAonIFO.png
                      Reply  Wed Mar 6 08:40:03 2013, Steve, Update, VAC, first scan of this pumpdown pd75m12d.pngpd75mRGA12d.png
                         Reply  Mon Jun 3 09:46:40 2013, Steve, Update, VAC, RGA scan at day 100 d100Mpd75.pngd100Mpd75p.pngbadguys.png
                            Reply  Mon Jul 15 14:12:15 2013, Steve, Update, VAC, RGA scan at day 144 pd75mRGA144d.pngpd75m144d.png
Message ID: 8206     Entry time: Fri Mar 1 14:50:36 2013     In reply to: 8155     Reply to this: 8238
Author: Steve 
Type: Update 
Category: VAC 
Subject: pumpdown day 8 

 Atm1, RGA background scan with TP3 pumping at day 7

Atm2,  IFO  CC1 = 1.1e-5 Torr at day 8

 Vacuum valve VM3 closed and VM2 opened, it switched the RGA to IFO scanning mode. CC4 = 3.5e-6 Torr

TP3 is pumping on the annuloses, their pressures are 1-3 mTorr

Vacuum condition reached the state of  Vacuum Normal. The MEDM screen "Current State" channel does not work because we disconnected some valves after the VME swap CAOS

Note: TP1 Maglev performance on Atm2  CC4 = 4e-9 Torr when VM2 valve was open

 

Attachment 1: bgTP3d7.png  120 kB  | Hide | Hide all | Show all
bgTP3d7.png
Attachment 2: RGAonIFO.png  151 kB  | Show | Hide all | Show all
ELOG V3.1.3-